OpenStudy (thatonegirl_):

Find the derivative of the function. y=sin(πx)^2

OpenStudy (irishboy123):

\(y=\sin(πx)^2\) OR \(y=\sin^2(πx)\)

OpenStudy (irishboy123):

well, do it then!

OpenStudy (thatonegirl_):

Jk the first one my b

OpenStudy (thatonegirl_):

Idk how lol

OpenStudy (thatonegirl_):

\[2\pi \cos \pi x\] this is what i got

OpenStudy (thatonegirl_):

but the answer is \[2\pi^2 x \cos(\pi x)^2\]

OpenStudy (irishboy123):

chain rule, is that what you did?

OpenStudy (s4sensitiveandshy):

\[\sin^2(\pi x) = \sin(\pi x)^2\]

OpenStudy (irishboy123):

\(\dfrac{d}{dx} \sin(πx)^2\) \(= \cos(πx)^2 \dfrac{d ((πx)^2}{dx} \) \(= \cos(πx)^2 2 (\pi x) \dfrac{d (πx) }{dx} \) \(= \cos(πx)^2 2 (\pi x) π\) \(= 2 \pi^2 x \cos(πx)^2 \) Chain Rule

OpenStudy (s4sensitiveandshy):

hmm. \[\sin( \pi x)^2 = \sin^2(\pi x) = (\sin(\pi x))^2\]

OpenStudy (thatonegirl_):

Whoa, are you supposed to use the power rule for this..??

OpenStudy (thatonegirl_):

because how i started it was \[y'=2\cos(\pi x) \]

OpenStudy (thatonegirl_):

then i used the chain rule to multiply it by π

OpenStudy (irishboy123):

do we agree that \(y=\sin(πx)^2\) \(=\sin u \) where \(u = (πx)^2 \)

OpenStudy (thatonegirl_):

Yea

OpenStudy (irishboy123):

so \(\dfrac{dy}{du} = cos u\)

OpenStudy (s4sensitiveandshy):

hmm

OpenStudy (thatonegirl_):

okay so when you take the derivative of some type of trig function and even if it's raised to a power, you just take the derivative of the trig and don't do anything with the power?

OpenStudy (irishboy123):

not sure what you mean but.... \(\dfrac{d}{dz} \sin z = \cos z\) \(\dfrac{d}{dz} \sin^2 z = 2 \sin z \cos z\) \(\dfrac{d}{dz} \cos^4 z = 4 \cos^3 z \dfrac{d}{dz} (\cos z) = - 4 \cos^3 z \sin z\)

OpenStudy (s4sensitiveandshy):

it is \[( \sin(\pi x) )^2\] apply the power rule alon the chain rule \[2(\sin(\pi x)) \ \cdot d' \sin(\pi x) \cdot d' \pi x\]

OpenStudy (s4sensitiveandshy):

\[y= (\sin(\pi x ))^2\] \[\ y'= 2 ( \sin(\pi x) 0* D~of~ \sin(\pi x) * D~of \pi x\] \[y' =2 \sin(\pi x) * \cos (\pi x) * \pi \]

OpenStudy (s4sensitiveandshy):

ignore 0

OpenStudy (irishboy123):

lol!! try reading the thread Shy & Sensitive!!

OpenStudy (thatonegirl_):

okay so the only thing im confused on is if sin(πx)^2 would be 2cos(πx)(π) or just cos(πx)(π) ??

OpenStudy (irishboy123):

it really depends upon what you are looking at!

OpenStudy (s4sensitiveandshy):

read what ?

OpenStudy (thatonegirl_):

For this problem what would it be?

OpenStudy (irishboy123):

if \(y = sin (\pi x)^2\) then we can define \(v = (\pi x)^2\) so \(y = \sin v\) and \(\dfrac{dy}{dv} = \cos v\) and \(\dfrac{dy}{dx} = \dfrac{dy}{dv}\bullet \dfrac{dv}{dx}\) and you can run it from there

OpenStudy (thatonegirl_):

so.. cos(πx)^2(π) ?

OpenStudy (irishboy123):

going tooooooo fast if \(v = (\pi x)^2\) then \(\dfrac{dv}{dx} = 2(\pi x) \pi = 2 \pi^2 x\)

OpenStudy (thatonegirl_):

okay i got that part but where does tht come into play in the original equation??

OpenStudy (irishboy123):

you gotta put it all together

OpenStudy (thatonegirl_):

cosv=cos2π^2x ?

OpenStudy (thatonegirl_):

waitt.. its \[cosv=2\pi^2xcos(\pi x) ^2\]

OpenStudy (thatonegirl_):

i think

OpenStudy (irishboy123):

but you're not sure!!

OpenStudy (thatonegirl_):

not completely but it sorta makes sense

OpenStudy (s4sensitiveandshy):

o.o

OpenStudy (irishboy123):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @IrishBoy123 it's a real bιτch, for sure \(\color{#0cbb34}{\text{End of Quote}}\)

OpenStudy (thatonegirl_):

lol xD yes i agree

OpenStudy (thatonegirl_):

I just don't get why you would take the derivative of the inside and multiply it separately if that makes sense

OpenStudy (s4sensitiveandshy):

:/

OpenStudy (irishboy123):

a = b^3 b = c^5 what is \(\dfrac {da}{dc}\) ?!!??

OpenStudy (thatonegirl_):

what is that even asking lol derivative of a?

OpenStudy (irishboy123):

lol!!!

OpenStudy (thatonegirl_):

:(((

OpenStudy (thatonegirl_):

ok imma forget about this one i'll figure it out tomorrow. I have another one to do lol

OpenStudy (s4sensitiveandshy):

@Thatonegirl_ \[y= \sin( \pi x )^2\] \[y= (\sin( \pi x ))^2\] power rule + chain rule \[y=2(\sin( \pi x)) * \cos(\pi x) * \pi\] not difficult

OpenStudy (s4sensitiveandshy):

i don't understand...how can we let v=(pix)^2 it is ( sin(pix)) squared

OpenStudy (s4sensitiveandshy):

\[y=x^n\]\[y'=nx^{n-1}\] \[y=(\sin (\pi x) )^2\] \[y'=2(\sin(\pi x))^{2-1} * D(\sin(\pi x)) * D(\pi x)\] D= derivative

OpenStudy (irishboy123):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @IrishBoy123 lol!! try reading the thread Shy & Sensitive!! \(\color{#0cbb34}{\text{End of Quote}}\)

OpenStudy (s4sensitiveandshy):

Originally Posted by @owl read what ? End of Quote: answer my Question.

OpenStudy (thatonegirl_):

I'll let you guys know how it goes tomorrow

OpenStudy (s4sensitiveandshy):

sure :)